OpenStudy (aglna):

___ 1.28 what is a rational number equivalent to this

OpenStudy (sshayer):

1.282828... =1+0.28+0.0028+0.000028+... =1+28(0.01+0.0001+0.000001+...) \[=1+28\left( \frac{ 1 }{ 100 }+\frac{ 1 }{ 10000 }+\frac{ 1 }{ 1000000 }+...\right)\] find r then\[s _{\infty}=\frac{ a }{ 1-r }\] then simplify

OpenStudy (sshayer):

or let x=1.282828... 100x=128.282828.. 100x-x=128 99x=128 x=?

OpenStudy (sshayer):

sorry correction 99x=127 x=?